OpenStudy (anonymous):

Let f(x) = x + 2 and g(x) = x^2 – 6x + 3. Find g(f(x)). a. x^2 – 6x + 5 b. x^2 – 5x + 5 c. x^2 – 2x – 5 d. x^2 + 10x + 19

OpenStudy (anonymous):

i think its a, am i right?

OpenStudy (ash2326):

f(x)=x+2 g(x)=\(x^2-6x+3\) g(f(x))=f(x)^2-6 f(x)+3 g(f(x))=(x+2)^2-6(x+2)+3 = x^2+4+4x-6x-12+3 =x^2-2x-5 option c

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so c?

OpenStudy (ash2326):

yeah

OpenStudy (anonymous):

ok thank u